OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in simplifying the following expression. Be sure to list restrictions. [Part 2 (2 points)] Use complete sentences to explain your process.

OpenStudy (anonymous):

\[(2x ^{2}+8x-42)\div(x ^{2}+12x+35)\]

OpenStudy (akashdeepdeb):

All you need to do here is factorize the numerator and the denominator! :) Split the middle term use FOIL! :D

OpenStudy (anonymous):

@satellite73 @Hero @nincompoop

hero (hero):

1. Factor the numerator and the denominator \[\frac{2(x + 7)(x - 3)}{(x+7)(x + 5)}\] 2. Cancel factors of one: \[\frac{2\cancel{(x + 7)}(x - 3)}{\cancel{(x+7)}(x + 5)}\] 3. Then write the simplified result \[\frac{2(x-3)}{x+5}\]

hero (hero):

Don't forget to list the restrictions.